OpenStudy (anonymous):

Little Miss Buffet takes all the money from her piggy bank and puts it into a savings account at her local bank. The bank promises an annual interest rate of 2.5% on the balance, compounded semiannually. How much will she have after one year if her initial deposit was $200? A. $205.00 B. $205.01 C. $205.02 D.$205.03

OpenStudy (anonymous):

what is 2.5% of 200$

OpenStudy (anonymous):

\[\frac{ 2.5 }{ 100 } \times 200=\]

OpenStudy (anonymous):

what would u get?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

good what is 200+5=

OpenStudy (anonymous):

205

OpenStudy (anonymous):

see it says "compounded semiannually" so in one year she ll get 205$

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

wait i am not sure i am right

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oright sorry about plz discard my previous answer its wrong i forgot there was a formula for this

OpenStudy (anonymous):

\[A=p(1+\frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

p is the initial amount which is 200 in our case r is the interest rate which is 0.025 always write in decimal form n is the number of compounding a year which is 2 because every 6 month t is the year which is 1

OpenStudy (anonymous):

plz sub values in formula and u ll get a correct answer :)

OpenStudy (anonymous):

\[A=200(1+\frac{ 0.025 }{ 2 })^{2}\]

OpenStudy (anonymous):

what u get?

OpenStudy (anonymous):

Do I have to solve for A?

OpenStudy (anonymous):

The answer is 205.03

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

answer is 205.03